A pendulum is executing simple harmoinic motion with a time period T if the length of pendulum is decreased to 36% the percentage chage in the time period of the pendulum is 
1)60%
2)40%
3)10%
4)20%

Grade:12

1 Answers

Arun
6 years ago
By formula T_{1}=2\pi \sqrt{\frac{l}{g}}\   
 
now length is decreased by 21% therefore new length = L – 36L/100 = 64 L/100          put in eq we get
T2 = 2\pi\sqrt36L/100 = T1 * 6/10     
 
T1 - T2 = 0.4 T1
so percentage change = \frac{change in time period }{original time period}\times 100   = 0.4*100   = 40%